# Efficient Neural Combinatorial Optimization Solver for the Min-max Heterogeneous Capacitated Vehicle Routing Problem

### v2

@@ Abstract
Numerous Neural Combinatorial Optimization (NCO) solvers have been proposed to address Vehicle Routing Problems (VRPs). However, most of these solvers focus exclusively on single-vehicle VRP variants, overlooking the more realistic min-max Heterogeneous Capacitated Vehicle Routing Problem (MMHCVRP), which involves multiple vehicles. Existing MMHCVRP solvers typically select a vehicle and its next node to visit at each decoding step, but often make myopic decoding decisions and overlook key properties of MMHCVRP, including local topological relationships, vehicle permutation invariance, and node symmetry, resulting in suboptimal performance. To better address these limitations, we propose ECHO, an efficient NCO solver. First, ECHO exploits the proposed dual-modality node encoder to capture local topological relationships among nodes. Subsequently, to mitigate myopic decisions, ECHO employs the proposed Parameter-Free Cross-Attention mechanism to prioritize the vehicle selected in the preceding decoding step. Finally, leveraging vehicle permutation invariance and node symmetry, we introduce a tailored data augment strategy for MMHCVRP to stabilize the Reinforcement Learning training process. To assess the performance of ECHO, we conduct extensive experiments. The experimental results demonstrate that ECHO outperforms state-of-the-art NCO solvers across varying numbers of vehicles and nodes, and exhibits well-performing generalization across both scales and distribution patterns. Finally, ablation studies validate the effectiveness of all proposed methods.
@@
{highlights}
We propose a dual-modality node encoder that jointly captures edge and node features.
We explicitly emphasize the importance of the previously selected vehicle in guiding the current decoding decision.
We propose a parameter-free cross attention mechanism to effectively incorporate features of the previously selected vehicle.
We develop an efficient data augment method tailored for MMHCVRP with complex constraints.
The proposed methods outperforms state-of-the-art methods across all evaluated settings.
@@ 1 Introduction
Vehicle Routing Problem (VRP) is a fundamental class of Combinatorial Optimization Problem (COP) with widespread applications in diverse domains, e.g., communication and transportation [airport_tits, bengio_machine_2021, xiao2023reinforcement, wu2025, 9115005]. Conventional algorithms for solving VRPs can be broadly divided into three categories, namely exact, approximation, and heuristic algorithms [li_research_2021, wu_incorporating_2023, SL_tits, ugas, wu_neural_2023]. However, these conventional algorithms cannot derive insights from historical VRP instances and lack the capability to solve multiple VRP instances in parallel, thus leading to significant computing overhead [wu_2024_survey].
@@
Recently, numerous Neural Combinatorial Optimization (NCO) solvers have been proposed to address the VRP [xiao2024improving, 2019_tits, 9901466, 10972166, xiao_distilling_2024]. These solvers can derive insights from historical instances and exploit GPU parallelism to process multiple VRP instances concurrently, achieving high-level performance with substantially reduced computational time [xin_multi-decoder_2021, glop]. However, most existing NCO solvers focus on either the unconstrained Traveling Salesman Problem (TSP) or the Capacitated Vehicle Routing Problem (CVRP) with only basic capacity constraints [mingzhao, mvmoe]. In practical applications, VRP variants often involve more complex attributes or constraints, thereby hindering the deployment of existing NCO research [li2024cada, drl]. For example, real-world scenarios frequently feature vehicles with heterogeneous attributes (e.g., capacity and speed), and fleet operators commonly aim to minimize the longest travel time among all vehicles (min-max) rather than the total time traveled by the entire fleet (min-sum). This problem is known as the min-max Heterogeneous Capacitated Vehicle Routing Problem (MMHCVRP).
Figure 1: Illustration on the decoding processes of 2D-Ptr [2dptr] (a) and the proposed ECHO (b). 2D-Ptr myopically relies on cumulative travel time to select the vehicle-node pair at each time step, which may lead to suboptimal solutions. Whereas the proposed ECHO alleviates this problem by emphasizing the vehicle selected in the preceding time step.
@@
Existing NCO methods for the MMHCVRP are typically trained via Reinforcement Learning (RL) and can be broadly categorized into two categories, namely AutoRegressive (AR) and Parallel-Autoregressive (PAR) solvers. AR-based solvers adopt an encoder-decoder architecture, where vehicle and node encoders embed vehicle and node features, respectively, and the decoder sequentially selects a vehicle and its next node to visit at each time step [drl, 2dptr]. PAR-based solvers similarly employ the node and vehicle encoders to embed their respective features. However, during decoding, each vehicle simultaneously selects a node to visit at each time step. This parallel selection enables PAR-based solvers to perform inference faster than their AR-based counterparts. However, the resulting high‑dimensional joint action space complicates training and may produce suboptimal solutions [dpn]. Moreover, conflicts arise when multiple vehicles simultaneously select the same node within a single time step. AR-based solvers also exhibit certain limitations. As shown in Fig. 1(a), these solvers disregard the priority of the vehicle selected in the preceding time step, relying solely on the accumulated travel time of each vehicle for decision-making. Consequently, at the third time step, vehicle $m_{3}$ is chosen to visit the node $n_{2}$. While this short‑sighted strategy may produce plausible solutions initially, it ultimately leads to suboptimal performance. Moreover, these solvers neglect key properties of the MMHCVRP, namely local topological relationships, vehicle permutation invariance, and node symmetry, which ultimately degrades their performance.
@@
To better address these limitations, we propose an Efficient Neural Combinatorial Optimization Solver for the Min-max Heterogeneous Capacitated Vehicle Routing Problem (ECHO). Specifically, to better capture local topological features, we replace the vanilla node encoder with our dual-modality node encoder, which employs a cross‑attention mechanism to integrate node and edge features (see Section 4.1 for more details). This design endows ECHO with robust generalization capabilities across both scales and distribution patterns. In addition, to mitigate myopic decisions, we propose a Parameter-Free Cross-Attention (PFCA) that prioritizes the node selected at the preceding time step during decoding. Specifically, PFCA efficiently integrates information about the vehicle selected in the t1t-1th time step into the node embeddings. Subsequently, ECHO selects a feasible vehicle-node pair at the $t$th time step based on the updated node embeddings and current vehicle embeddings (see Section 4.2 for more details). As shown in Fig. 1(b), in the third decoding step, vehicle $m_{2}$ is selected to visit the node $n_{2}$ instead of vehicle $m_{3}$, ultimately yielding a superior solution. Finally, we propose a tailored data augment method for MMHCVRP, which effectively generates diverse instances by leveraging vehicle permutation invariance and node symmetry to stabilize the RL training process.
@@
To assess the effectiveness of the proposed ECHO, we conduct extensive experiments on MMHCVRP instances with different numbers of vehicles and nodes. The experimental results demonstrate that ECHO outperforms the other NCO solvers designed for MMHCVRP, including two AR-based NCO solvers, namely DRL [drl] and 2D-Ptr [2dptr], as well as one PAR-based solver, namely PARCO [parco]. Notably, ECHO reduces the average gap by approximately 3% compared to the state-of-the-art (SOTA) solver PARCO across all vehicle and node scales (see Section 5.2). In addition, ECHO exhibits robust cross-scale and cross-distribution generalization ability. Finally, ablation studies demonstrate the effectiveness of the proposed dual-modality node encoder, PFCA mechanism, and data augment method.
@@ 2.1 NCO solvers for VRP
Existing NCO solvers for VRP can be broadly categorized into two types, namely Learning to Construct and Learning to Improve solvers [li2025destroy, wu_learning_2022, 10259654]. Specifically, inspired by sequence‑to‑sequence models in the machine translation task, Learning to Construct solvers sequentially select unvisited nodes and append them to partial solutions to construct the complete solutions [vinyals_pointer_2015, kool_attention_2019, luo_neural_2023]. In contrast, Learning to Improve solvers, motivated by ruin‑and‑repair heuristics, iteratively perturb and repair current complete solutions to explore (sub-)optimal solutions within a given time frame [ma_learning_2021, ma_learning_2023, chen_learning_2019]. Among these two paradigms, the Learning‑to‑Construct solver is more prevalent.
As highlighted in the recent survey by Wu et al. [wu_2024_survey], although these solvers have achieved significant progress, they still encounter certain limitations. Specifically, existing NCO solvers exhibit poor cross‑scale and cross-distribution generalization abilities and cannot solve large-scale VRPs in real-time [xiao2025, glop, luo2025boosting, dgl]. In addition, their applicability remains largely confined to TSP and CVRP, with limited success on more complex VRP variants [mvmoe, li2024cada]. An important variant is the min-max VRP, where the objective is to minimize the maximum route duration across the entire vehicle fleet. Although several studies have examined the min-max TSP [Liang_Splitnet_2023, 10040987, kim2023learning, dpn, et], they overlook vehicle heterogeneity, and the simplified constraints of min-max TSP hinder applicability to real-world routing scenarios. Consequently, it is essential to investigate the MMHCVRP, which incorporates these additional practical constraints.
@@ 2.2 NCO solvers Specifically Designed for MMHCVRP
To better solve the more challenging MMHCVRP, Li et al. [drl] introduced DRL, the first AR-based solver tailored for this problem. Specifically, DRL employs an attention-based encoder and a Fully Connected Network (FCN) to embed vehicle and node features, respectively. During decoding, at each time step, it employs a policy network that first selects a vehicle and then chooses that vehicle’s next node to visit, iteratively constructing complete routes until all customer nodes have been visited. Building on the prior study [drl], Liu et al. [2dptr] proposed 2D-Ptr, which incorporates a more efficient decoder. Specifically, at each time step, 2D-Ptr directly computes the dot product between vehicle and node embeddings, and selects the vehicle-node pair with the highest scores as the current vehicle and its next node to visit. Compared to DRL’s sequential first-vehicle‑then‑node selection, 2D‑Ptr’s joint selection of vehicle-node pair significantly enhances the computational efficiency. However, both AR-based NCO solvers overlook critical MMHCVRP properties, namely local topological relationships, vehicle permutation invariance, and node symmetry. In addition, they disregard the priority of the previously selected vehicle during the decoding process, resulting in suboptimal solutions.
Unlike AR methods that sequentially select one vehicle and its next node to visit at each time step, Berto et al. [parco] proposed PARCO, a PAR-based solver for MMHCVRP that simultaneously selects each vehicle’s next node at every decoding step. Theoretically, PAR‑based solvers require only $\frac{1}{M}$ of the time steps of AR‑based solvers, where $M$ denotes the number of vehicles, thereby achieving superior efficiency. However, PAR‑based solvers suffer from a high‑dimensional joint action space that complicates training and necessitates conflict handlers among vehicles, thereby limiting their performance [dpn].
@@ 3.1 Definition of MMHCVRP
Consider an MMHCVRP instance $G(\mathbb{M},\mathbb{N})$, where $\mathbb{M}=\{m_{i}\}_{i=1}^{M}$ denotes a fleet of $M$ heterogeneous vehicles, and $\mathbb{N}=\{n_{j}\}_{j=0}^{N}$ denotes a set of nodes comprising a depot $n_{0}$ and the number of $N$ customers. In this problem, each fully loaded vehicle departs from the depot and visits a sequence of customers to satisfy their demands, subject to the constraints that each customer is visited exactly once and the load carried by any vehicle along its route does not exceed its capacity. Each vehicle $m_{i}$ is characterized by two attributes, namely, capacity $\rho_{i}$ and speed $\chi_{i}$. Each node $n_{i}$ is described by three attributes, namely, coordinates $x_{j}$ and $y_{j}$, as well as demand $q_{j}$. In addition, the objective is to minimize the maximum route duration among all vehicles, which are defined as follows:
$\min\max_{i\in\mathbb{M}}(\sum_{j=1}^{|\gamma_{i}|}\frac{dist(\gamma_{i}^{j-1},\gamma_{i}^{j})}{\chi_{i}}+\frac{dist(\gamma_{i}^{|\gamma_{i}|},\gamma_{i}^{0})}{\chi_{i}}),$ (1)
@@ 3.2 Markov Decision Process
Following the prior studies [drl, 2dptr], we adopt the AR paradigm and define an identical MDP to derive the visitation sequence $\tau=\{(m^{t},n^{t}),t=\{0,1,...,T-1\}\}$ for solving the MMHCVRP. Specifically, the MDP comprises the following five components:
@@ 3.2.3 Transaction
Given the current state $s^{t}$ and action $a^{t}=(m_{i}^{t},n_{j}^{t})$, the next state $s^{t+1}=f(s^{t},a^{t})$ is obtained by updating each vehicle’s accumulated usage capacity, accumulated travel time, and current location, as well as each node’s demand. For every vehicle $k\in\{1,\cdots,M\}$ and node $l\in\{0,\cdots,N\}$, these updates are defined as follows:
@@ 3.2.4 Reward
The total reward $R$ is defined as the negative of the longest travel time across all vehicles, as follows:

### v3

@@ Abstract
Numerous Neural Combinatorial Optimization (NCO) solvers have been proposed to address Vehicle Routing Problems (VRPs). However, most of these solvers focus exclusively on single-vehicle VRP variants, overlooking the more realistic Min-Max Heterogeneous Capacitated Vehicle Routing Problem (MMHCVRP), which involves multiple vehicles. Existing MMHCVRP solvers typically select a vehicle and its next node to visit at each decoding step, but often make myopic decisions during decoding and overlook key properties of MMHCVRP, including local topological relationships, vehicle permutation invariance, and node symmetry, resulting in suboptimal performance. To better address these limitations, we propose ECHO, an efficient NCO solver. First, ECHO exploits the proposed dual-modality node encoder to capture local topological relationships among nodes. Subsequently, to mitigate myopic decisions, ECHO employs the proposed Parameter-Free Cross-Attention mechanism to prioritize the vehicle selected in the preceding decoding step. Finally, leveraging vehicle permutation invariance and node symmetry, we introduce a tailored data augment strategy for MMHCVRP to stabilize the Reinforcement Learning training process. To assess the performance of ECHO, we conduct extensive experiments. The experimental results demonstrate that ECHO outperforms state-of-the-art (SOTA) NCO solvers across varying numbers of vehicles and nodes, and exhibits well-performing generalization across both scales and distribution patterns. Finally, ablation studies validate the effectiveness of all proposed methods.
@@ 1 Introduction
Vehicle Routing Problem (VRP) is a fundamental class of Combinatorial Optimization Problem (COP) with widespread applications in diverse domains, e.g., communication and transportation (Wu et al., 2023b; Bengio et al., 2021; Xiao et al., 2025a; Wu et al., 2025; Yang et al., 2020). Conventional algorithms for solving VRPs can be broadly divided into three categories, namely exact, approximation, and heuristic algorithms (Li et al., 2021; Wu et al., 2023a; Song et al., 2025; Wu et al., 2024a). However, these conventional algorithms cannot derive insights from historical VRP instances and lack the capability to solve multiple VRP instances in parallel, thus leading to significant computing overhead (Wu et al., 2024b).
@@
Recently, numerous Neural Combinatorial Optimization (NCO) solvers have been proposed to address the VRP (Xiao et al., 2025b; Zhang et al., 2023; Li et al., 2024; Xiao et al., 2024; Luo and Li, 2025; Li et al., 2026). These solvers can derive insights from historical instances and exploit GPU parallelism to process multiple VRP instances concurrently, achieving high-level performance with substantially reduced computational time (Xin et al., 2021; Ye et al., 2024b). However, most existing NCO solvers focus on either the unconstrained Traveling Salesman Problem (TSP) or the Capacitated Vehicle Routing Problem (CVRP) with only basic capacity constraints (Wang et al., 2025a; Zhou et al., 2024). In practical applications, VRP variants often involve more complex attributes or constraints, thereby hindering the deployment of existing NCO research (Li et al., 2025a, 2022). For example, real-world scenarios frequently feature vehicles with heterogeneous attributes (e.g., capacity and speed), and fleet operators commonly aim to minimize the longest travel time among all vehicles (min-max) rather than the total time traveled by the entire fleet (min-sum). This problem is known as the Min-Max Heterogeneous Capacitated Vehicle Routing Problem (MMHCVRP).
@@
Existing NCO methods for the MMHCVRP are typically trained via Reinforcement Learning (RL) and can be broadly categorized into two categories, namely AutoRegressive (AR) and Parallel-Autoregressive (PAR) solvers. AR-based solvers adopt an encoder-decoder architecture, where vehicle and node encoders embed vehicle and node features, respectively, and the decoder sequentially selects a vehicle and its next node to visit at each time step (Li et al., 2022; Liu et al., 2024). PAR-based solvers similarly employ the node and vehicle encoders to embed their respective features. However, during decoding, each vehicle simultaneously selects a node to visit at each time step. This parallel selection enables PAR-based solvers to achieve faster inference than their AR-based counterparts. However, the resulting high‑dimensional joint action space complicates training and may produce suboptimal solutions (Zheng et al., 2024). Moreover, conflicts arise when multiple vehicles simultaneously select the same node within a single time step. AR-based solvers also exhibit certain limitations. As shown in Figure 3(a), the SOTA AR-based solver 2D-Ptr disregards the priority of the vehicle selected in the preceding time step and rely solely on the accumulated travel time of each vehicle for decision-making. Consequently, at the third time step, vehicle $m_{3}$ is chosen to visit the node $n_{2}$. While this short‑sighted strategy may produce plausible solutions initially, it ultimately leads to suboptimal performance. Moreover, both these AR-based and NAR-based solvers neglect key properties of the MMHCVRP, namely local topological relationships, vehicle permutation invariance, and node symmetry, which ultimately degrades their performance.
@@
To better address these limitations, we propose an Efficient Neural Combinatorial Optimization Solver for the Min-max Heterogeneous Capacitated Vehicle Routing Problem (ECHO). Specifically, to better capture local topological features, we replace the vanilla node encoder with our dual-modality node encoder, which employs a cross‑attention mechanism to integrate node and edge features (see Section 4.1 for more details). This design endows ECHO with robust generalization capabilities across both scales and distribution patterns. In addition, to mitigate myopic decisions, we propose a Parameter-Free Cross-Attention (PFCA) that prioritizes the node selected at the preceding time step during decoding. Specifically, PFCA efficiently integrates information about the vehicle selected in the t1t-1th time step into the node embeddings. Subsequently, ECHO selects a feasible vehicle-node pair at the $t$th time step based on the updated node embeddings and current vehicle embeddings (see Section 4.2 for more details). As shown in Figure 3(b), in the third decoding step, vehicle $m_{2}$ is selected to visit the node $n_{2}$ instead of vehicle $m_{3}$, ultimately yielding a superior solution. To the best of our knowledge, ECHO is the first work to explicitly model historically selected vehicle information and design a tailored mechanism to leverage it for guiding sequential decision-making, thereby effectively alleviating myopic decisions. Finally, we propose a tailored data augment method for MMHCVRP, which effectively generates diverse instances by leveraging vehicle permutation invariance and node symmetry to stabilize the RL training process.
@@
To assess the effectiveness of the proposed ECHO, we conduct extensive experiments on MMHCVRP instances with different numbers of vehicles and nodes. The experimental results demonstrate that ECHO outperforms the other NCO solvers designed for MMHCVRP, including two AR-based NCO solvers, namely DRL (Li et al., 2022) and 2D-Ptr (Liu et al., 2024), as well as one PAR-based solver, namely PARCO (Berto et al., 2025). Notably, ECHO reduces the average gap by approximately 3% compared to the SOTA solver PARCO across all vehicle and node scales (see Section 5.2). In addition, ECHO exhibits robust cross-scale and cross-distribution generalization ability. Finally, ablation studies demonstrate the effectiveness of the proposed dual-modality node encoder, PFCA mechanism, and data augment method, while the case study further illustrates the effectiveness of PFCA in avoiding the neglect of the priority of previously selected vehicles.
@@
The remainder of this paper is organized as follows: Section 2 reviews the related work in the field of NCO. Section 3 presents the definition of MMHCVRP and the MDP formula used in our method. Section 4 describes the architecture of ECHO and the proposed data augment strategy. Section 5 discusses the experimental results. Section 6 discusses the limitations of the proposed ECHO. Finally, Section 7 draws the conclusion.
@@ 2.1 NCO solvers for VRP
Existing NCO solvers for VRP can be broadly categorized into two types, namely Learning to Construct and Learning to Improve solvers (Li et al., 2025b; Wu et al., 2022; Guan et al., 2025). Inspired by sequence‑to‑sequence models in the machine translation task, Learning to Construct solvers sequentially select unvisited nodes and append them to partial solutions to construct the complete solutions (Vinyals et al., 2015; Kool et al., 2019; Luo et al., 2023). In contrast, Learning to Improve solvers, motivated by ruin‑and‑repair heuristics, iteratively perturb and repair current complete solutions to explore (sub-)optimal solutions within a given time frame (Ma et al., 2021, 2023; Chen and Tian, 2019). Among these two paradigms, the Learning‑to‑Construct solver is more prevalent.
As highlighted in the recent survey by Wu et al. (Wu et al., 2024b), although these solvers have achieved significant progress, they still encounter certain limitations. Specifically, existing NCO solvers exhibit poor cross‑scale and cross-distribution generalization abilities and cannot solve large-scale VRPs in real-time (Xiao et al., 2026; Ye et al., 2024b; Luo et al., 2025; Xiao et al., 2025c). In addition, their applicability remains largely confined to TSP and CVRP, with limited success on more complex VRP variants (Zhou et al., 2024; Li et al., 2025a). An important variant is the min-max VRP, where the objective is to minimize the maximum route duration across the entire vehicle fleet. Although several studies have examined the min-max TSP (Liang et al., 2023; Gao et al., 2024; Kim et al., 2023; Zheng et al., 2024; Son et al., 2024), they overlook vehicle heterogeneity, and the simplified constraints of min-max TSP hinder applicability to real-world routing scenarios. Consequently, it is essential to investigate the MMHCVRP, which incorporates these additional practical constraints.
@@ 2.2 NCO solvers Specifically Designed for MMHCVRP
To better solve the more challenging MMHCVRP, Li et al. (2022) introduced DRL, the first AR-based solver tailored for this problem. Specifically, DRL employs an attention-based encoder and a Fully Connected Network (FCN) to embed vehicle and node features, respectively. During decoding, at each time step, it employs a policy network that first selects a vehicle and then chooses that vehicle’s next node to visit, iteratively constructing complete routes until all customer nodes have been visited. Building on the prior study (Li et al., 2022), Liu et al. (2024) proposed 2D-Ptr, which incorporates a more efficient decoder. Specifically, at each time step, 2D-Ptr directly computes the dot product between vehicle and node embeddings, and selects the vehicle-node pair with the highest scores as the current vehicle and its next node to visit. Compared to DRL’s sequential first-vehicle‑then‑node selection, 2D‑Ptr’s joint selection of vehicle-node pair significantly enhances the computational efficiency. However, both AR-based NCO solvers overlook critical MMHCVRP properties, namely local topological relationships, vehicle permutation invariance, and node symmetry. In addition, they disregard the priority of the previously selected vehicle during the decoding process, resulting in suboptimal solutions. To better solve these limitations, we propose ECHO (see Section 4 for more details).
Unlike AR methods that sequentially select one vehicle and its next node to visit at each time step, Berto et al. (2025) proposed PARCO, a PAR-based solver for MMHCVRP that simultaneously selects each vehicle’s next node at every decoding step. Theoretically, PAR‑based solvers require only $\frac{1}{M}$ of the time steps of AR‑based solvers, where $M$ denotes the number of vehicles, thereby achieving superior efficiency. However, PARCO suffers from a high‑dimensional joint action space that complicates training and necessitates conflict handlers among vehicles. Consequently, although PARCO adopts a strong training paradigm, its performance remains constrained (Zheng et al., 2024). In contrast, we extend the augmentation scheme used in PARCO and introduce a more specifically tailored architecture, which jointly contributes to the superior performance of ECHO.
@@ 3.1 Definition of MMHCVRP
Consider an MMHCVRP instance $G(\mathbb{M},\mathbb{N})$, where $\mathbb{M}=\{m_{i}\}_{i=1}^{M}$ denotes a fleet of $M$ heterogeneous vehicles, and $\mathbb{N}=\{n_{j}\}_{j=0}^{N}$ denotes a set of nodes comprising a depot $n_{0}$ and the number of $N$ customers. In this problem, each fully loaded vehicle departs from the depot and visits a sequence of customers to satisfy their demands, subject to the constraints that each customer is visited exactly once and the load carried by any vehicle along its route does not exceed its capacity. Each vehicle $m_{i}$ is characterized by two attributes, namely, capacity $\rho_{i}$ and speed $\chi_{i}$. Each node $n_{i}$ is described by three attributes, namely, coordinates $x_{j}$ and $y_{j}$, as well as demand $q_{j}$. In addition, the objective is to minimize the maximum route duration among all vehicles, which are defined in Eq. (1).
$\min\max_{i\in\mathbb{M}}(\sum_{j=1}^{|\gamma_{i}|}\frac{dist(\gamma_{i}^{j-1},\gamma_{i}^{j})}{\chi_{i}}),$ (1)
@@ 3.2 Markov Decision Process
Following the prior studies (Li et al., 2022; Liu et al., 2024), we adopt the AR paradigm and define an identical MDP to derive the visitation sequence $\tau=\{(m^{t},n^{t}),t=\{0,1,...,T-1\}\}$ for solving the MMHCVRP. Specifically, the MDP comprises the following five components:
@@ 3.2.3 Transaction
[…]
@@ 3.2.4 Reward
The total reward $R$ is defined as the negative of the longest travel time across all vehicles, as given in Eq. (6).